# Partial Wavefunction Collapse Under Repeated Weak Measurement of a non-Conserved Observable

### v1

@@ Abstract
Two hallmarks of quantum non-demolition (QND) measurement are the ensemble-level conservation of the expectation value of the measured observable $A$ and the eventual, inevitable collapse of the system into some eigenstate of $A$. This requires that $A$ commutes with $H$, the system’s Hamiltonian. In what we term “non-QND” measurement, $A$ does not commute with $H$ and these two characteristics clearly cannot be present as the system’s dynamics prevent $\langle A\rangle$ from reaching a stable value. However, in this paper we find that under non-QND conditions, QND-like behavior can still arise, but is seen in the behavior of a secondary observable we call $B$, with the condition that $B$ commutes with both $A$ and $H$. In such cases, the expectation value of $B$ is conserved and the system at least partially collapses with respect to eigenstates of $B$. We show as an example how this surprising result applies to a Heisenberg chain, where we demonstrate that local measurements on a single site can reveal information about the entire system.
@@ I Introduction
Weak measurement is a type of measurement in which the system of interest is not directly projected in order to avoid immediate collapse. This can be achieved by entangling the system with an ancilla which is itself then projectively measured. In this way, controlling the parameters coupling the ancilla to the system results in a tunable trade-off between the amount of information obtained and the degree to which the original system state is disturbed (for an overview of weak measurement, see [1]). When the observable to be measured commutes with the system Hamiltonian this produces a quantum non-demolition (QND) measurement [2]. Importantly, non-demolition does not mean that the system is unchanged, just that the ensemble average of the expectation value of the measured observable is unchanged by the measurement. It has been shown that repeated QND weak measurements converge to an eigenstate of the measured observable, asymptotically reproducing a projective measurement and the associated wavefunction collapse [3]. These two properties have made QND weak measurement an appealing tool for experiment, as it behaves in a way that we expect measurements to behave, while giving the experimenter more control than can be achieved with only projective measurements.
@@
QND measurement has received significant attention over the years, particularly in the context of experiment. Early work on foundations and applications of QND measurements can be seen in [2] and [4]. More recently, QND weak measurement was used in [5] in a breakthrough experiment that was able to intercept and reverse individual quantum jumps as they happened. QND has also featured heavily in the context of understanding wavefunction collapse and the “measurement problem”. See, for example, [6, 7]. Collapse under weak measurement was also examined by Korotkov in [8] in the context of continuous measurement of a double quantum dot.
@@
In contrast to QND measurement, one might instead consider a case in which the measured operator does not commute with the system Hamiltonian. Here we call this situation non-QND measurement. In QND the measurement back-action accumulates uninterrupted by any system dynamics, culminating in eventual collapse to an eigenstate of the measured operator as seen in fig. 1, yet in non-QND sequential measurement the system dynamics do interfere with the back-action, clearly rendering it impossible for collapse to an eigenstate of the measured operator to occur. So, for the behavior of a system subjected to this type of measurement, we would perhaps expect something resembling the early-time behavior of the QND case, but with no eventual collapse. One would therefore expect that subjecting a system to such a measurement protocol would not reveal any useful information. In [8] a non-QND case was examined, however the conditions considered were too restricted for our main result to have been noticed in that context.
@@
In this paper we theoretically examine the consequences of non-QND measurement processes and present as an example a numerical simulation of such a process being applied to a particular system, the Heisenberg chain. Our results indicate that the naïve expectation of non-collapse does not fully reflect reality, as a non-QND measurement process does indeed extract information from the system and lead to a partial ”collapse” of the system state. In the particular case of a Heisenberg chain, we find that a particular measurement protocol causes the z-component of the spin on the end of the chain to oscillate with sharp Fourier peaks that correspond to energy gaps in the spectrum of the chain. Further, we show that these oscillations arise as a result of the chain collapsing into a sector with a definite total z-component. We demonstrate that such collapse can be expected under fairly general circumstances for any operator which commutes with both the system Hamiltonian and the measured operator. This indicates that there is the possibility of gaining useful information from non-QND sequential measurement protocols.
@@ II Repeated Weak Measurement
Defining $\tau$ to be the (dimensionless) time over which the pre-measurement is performed, the joint system-ancilla density matrix $\alpha$ becomes
@@
For completeness and since it will be useful later, the expression for $\operatorname{Tr}_{\mu}(\operatorname{\mathbb{P}}_{\pm}\alpha_{k}(\tau))$ to second order is
$\operatorname{Tr}_{\mu}(\operatorname{\mathbb{P}}_{\pm}\alpha_{k}(\tau))$ $=\frac{\rho_{k}}{2}+\frac{i\tau}{2}[\rho_{k},H_{0}]+\frac{\tau^{2}}{2}(H_{0}\rho_{k}H_{0}-\frac{1}{2}\{H_{0}^{2},\rho_{k}\})\pm\gamma\tau(\rho_{k}A+A\rho_{k})$ (7)
@@
II.1 Expectation Value of Secondary Observable
@@
We claim that each $v_{nm}$ is a bounded super-martingale, and thus by the martingale convergence theorem each one converges almost surely [10]. That is, after some finite (but unknown) time it is guaranteed that each $v_{nm}$ will have converged to some stable value. We will now support the claim that the $v_{nm}$ are super-martingales, and give a condition for when they all converge to 00. It is evident that $v(\rho)$ is positive semi-definite (showing boundedness), and inspection reveals that $v(\rho)=0$ if and only if $\rho$ is collapsed into an eigensector of $B$. Thus, if we can show that $\mathbb{E}(v(\rho_{k+1})-v(\rho_{k}))\leq 0$ then $v(\rho)$ is a super-martingale and its convergence will tell us about collapse w.r.t. $B$.
@@
$\mathbb{E}(\Delta v_{mn})=-\frac{2\gamma^{2}\tau^{2}}{v_{mn}(\rho)^{3}}\left(\operatorname{Tr}(\operatorname{\mathbb{P}}_{n}\rho_{k})\operatorname{Tr}(\operatorname{\mathbb{P}}_{m}A\rho_{k})-\operatorname{Tr}(\operatorname{\mathbb{P}}_{m}\rho_{k})\operatorname{Tr}(\operatorname{\mathbb{P}}_{n}A\rho_{k})\right)^{2}.$ (11)
Since the term in parentheses is squared and we already know $v_{nm}\geq 0$, it is clear that $\mathbb{E}(\Delta v)\leq 0$ and so $v$ is a super-martingale which converges almost surely. There are then two remaining possibilities concerning the limit of $v$: either $v$ converges to 0 signifying that the state converges into a single eigensector of $B$, or $v$ converges to some other value. The latter case requires that an equilibrium be reached in which the term in parentheses is 00 at all times. Manipulating this term, we can rephrase this condition as being that the convergence stops at a step labelled $k_{0}$ when the condition
@@
is satisfied for all $n,m$ with $k\geq k_{0}$. Physically, this condition is that the expectation value of $A$ is at all times independent of any post-selection of an eigenvalue of $B$. To make contact with the QND case we again set $B=A$, giving the condition that $\langle A\rangle$ must not depend on any post-selection on $A$. This is clearly only met when the state occupies just one eigensector of $A$. So, we recover the QND result of guaranteed convergence to an eigenvalue.
@@
Considering this condition more closely, we can see a direct link between information extraction and collapse. As seen in Eq. 5, the information extracted by the weak measurement is information about $\langle A\rangle$. Rephrasing the collapse-halting condition, we can say that the collapse stops when information about $\langle A\rangle$ ceases to be correlated with the value of $B$. So, we see that the weak measurement process indirectly extracts information about $B$, and that as long as it does so the system slowly collapses towards a $B$-eigensector. The collapse then stops either when an eigensector is reached and there is no additional information about $B$ to be gained, or when eigensectors of $B$ can no longer be distinguished via $\langle A\rangle$ and so the weak measurement stops extracting information about $B$.
@@
It is of further interest to examine what is required for a system to collapse only partially with respect to $B$. The non-collapse condition can be rephrased as requiring a state in which $\langle A\rangle$ is identical in two or more $B$-sectors for all $k>k_{0}$. This necessarily requires $\langle A\rangle$ in each of those sectors to have the same time-dependence to maintain the balance. There are two main sources of time-dependence present: the unitary evolution under which phases are introduced to energy eigenstates according to their eigenvalue, and the stochastic evolution introduced by the repeated measurements which adjust the weights of different $A$ eigenstates. Assuming that these two sources of time-dependence affect $\langle A\rangle$ independently over a single measurement cycle since one is deterministic and the other stochastic, this means that both sources of time-dependence should separately be identical across the involved sectors. Since the unitary evolution acts on energy eigenstates, this means that the hypothetical non-collapsing state should in each sector have an equivalent decomposition into energy eigenstates so that the unitary evolution acts equivalently on all sectors. Similarly, since the stochastic evolution acts on $A$ eigenstates, an analogous condition should hold for the composition of the state across sectors in terms of $A$ eigenstates. Thus, in order to be able to construct a non-collapsing state, it must be possible to construct a state which has in two or more sectors equivalent decompositions with respect to both energy eigenstates and $A$ eigenstates. Further, we should expect these decompositions to have to be identical on a per-component basis, i.e. each involved energy eigenstate in the relevant sectors should have the same decomposition into $A$ eigenstates within the sector. Otherwise, the stochastic evolution would unequally affect the weights of that energy eigenstate in different sectors, disrupting the balance of having identical unitary time-dependence. We have thus far not found any nontrivial examples of such a condition being met.
@@ III Application to Heisenberg Chain
Refer to caption(a)AmplitudeFrequency ($J/2\pi$)
@@
Figure 2: Behavior of the 4-site Heisenberg chain subjected to repeated weak measurement of $\sigma^{z}_{0}$. a) Fourier Transform of $\langle\sigma^{z}_{0}\rangle(t)$ showing sharp peaks corresponding to certain energy gaps in the spectrum of the chain. b) Overlap of the chain’s state with sectors of total z-component shows a collapse into one sector following stochastic evolution. c) Distribution of expectation values of total z-component for a sample of 1,000 chains shown after certain numbers of measurements. It is clear that each individual run collapses into an eigensector of total z-component, while the ensemble average expectation value remains 0, demonstrating QND-like behavior.
@@
$H=\sum_{ij}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}$ (13)
We choose to label the spins along the chain from 00 to $N$ so that e.g. $\vec{\sigma}_{0}$ refers to the first spin on the chain. The operator we weakly measure is $\sigma^{z}_{0}$, the z-component of the first spin. To make the connection clear, this $\sigma^{z}_{0}$ operator corresponds to $A$ in the previous section. The operator corresponding to $B$ will be the total z-component of the chain, which we refer to here as $m_{s}$. It is easy to verify that the required condition $[A,B]=[B,H]=0$ is met by this set of operators. The experimentally accessible information from this protocol will be the weak measurement record of $+1$ and $-1$ outcomes of the ancilla, but here we also explicitly track the system’s state over time in order to gain more insight into how it behaves.
@@
Of additional interest is the Fourier transform of $\langle\sigma^{z}_{0}\rangle$, which reveals oscillatory behavior in the z-component of the measured spin with several sharp peaks in frequency space. Examination of these peaks reveals that their spacing corresponds to certain spacings in the energy spectrum of the Heisenberg chain. In particular, the result shown in the figure represents spacings between $m_{s}=1$ eigenstates according to the selection rule $\Delta S=0,1$. That is, pairs of eigenstates with $m_{s}=1$ for both states and a difference of 0 or 1 in total spin contribute to the oscillation of the end spin. This finding of frequency peaks which obey this particular selection rule is specific to the Heisenberg chain but supports the finding of collapse into a single $m_{s}$ sector, as otherwise additional peaks would be present. These oscillations would also be seen in the experimentally accessible measurement record, though heavily suppressed by noise. As such, our findings can be experimentally tested as long as a sufficient quantity of data is collected to overcome the statistical penalty of weak measurement.

### v2

@@ Abstract
Two hallmarks of quantum non-demolition (QND) measurement are the ensemble-level conservation of the expectation value of the measured observable $A$ and the eventual, inevitable collapse of the system into some eigenstate of $A$. This requires that $A$ commutes with $H$, the system’s Hamiltonian. In what we term “Auxiliary Observable QND” measurement, $A$ does not commute with $H$ and the above two characteristics clearly cannot be present as the system’s dynamics prevent $\langle A\rangle$ from reaching a definite value. However, in this paper we find that under such a measurement QND behavior still arises, but is seen in the behavior of a secondary “target” observable we call $B$, with the condition that $B$ commutes with both $A$ and $H$. In such cases, the expectation value of $B$ is conserved and the system at least partially collapses with respect to eigenstates of $B$. We show as an example how this surprising result applies to a Heisenberg chain, where we demonstrate that local measurements on a single site can reveal information about the spectrum of an entire system, a finding which may be of practical use in experiments.
@@ I Introduction
An important concept in the theory of quantum measurements is the notion of wavefunction collapse, where following a projective measurement of a system prepared in a superposition of eigenstates, any subsequent measurements of the same observable on the same system yield the same eigenvalue as the first measurement. Weak measurement is a type of measurement in which the system of interest is not directly projected in order to avoid this immediate collapse. This can be achieved by entangling the system with an ancillary degree of freedom which is itself then projectively measured. In this way, controlling the parameters coupling the ancilla to the system results in a tunable trade-off between the amount of information obtained and the degree to which the original system state is disturbed (for an overview of weak measurement, see [1]). When the observable to be measured commutes with the system Hamiltonian this produces a quantum non-demolition (QND) measurement [2]. Importantly, non-demolition does not mean that the system is unchanged, just that the ensemble average of the expectation value of the measured observable is unchanged by the measurement. It has been shown that repeated QND weak measurements converge to an eigenstate of the measured observable, asymptotically reproducing a projective measurement and the associated wavefunction collapse [3]. In Fig. 1, we illustrate this behavior in the example of QND weak measurement of a single qubit. In particular, panel (c) highlights the two important properties of eventual collapse and conservation of ensemble-average expectation value. These two properties have made QND weak measurement an appealing tool for experiment, as it behaves in a way that one expects measurements to behave, while giving the experimenter more control than can be achieved with only projective measurements.
@@
QND measurement has received significant attention over the years, particularly in the context of experiment. Early work on foundations and applications of QND measurements can be seen in [2] and [4]. More recently, QND weak measurement was used in a breakthrough experiment that was able to intercept and reverse individual quantum jumps as they happened [5]. QND has also featured heavily in the context of understanding wavefunction collapse and the “measurement problem” [6, 7]. Collapse under weak measurement has also been examined by Korotkov in the context of continuous measurement of a double quantum dot [8].
@@
In contrast to QND measurement, one might instead consider a case in which the measured operator does not commute with the system Hamiltonian. For reasons that will soon become clear, we call this situation Auxiliary Observable QND (AO-QND) measurement. In ordinary QND the measurement back-action accumulates uninterrupted by any system dynamics, culminating in eventual collapse to an eigenstate of the measured operator, yet in AO-QND sequential measurement the system dynamics do interfere with the back-action, clearly rendering it impossible for collapse to an eigenstate of the measured operator to occur. So, for the behavior of a system subjected to this type of measurement, we would perhaps expect something resembling the early-time behavior of the QND case, but with no eventual collapse. One would therefore expect that subjecting a system to such a measurement protocol would not reveal any useful information. Indeed, such a protocol has been determined to behave as if the measurement basis was chosen randomly for each measurement, making gleaning useful information a challenging task, although it has also been shown that the protocol could at least be used for quantum state tomography [9].
@@
In this paper we examine the consequences of AO-QND measurement processes and present as an example a numerical simulation of such a process being applied to a particular system, the Heisenberg chain. Our results indicate that the naïve expectation of non-collapse does not fully reflect reality, as an AO-QND measurement process does indeed extract information from the system and lead to at least a partial collapse of the system state. In section II we approach the problem theoretically and prove that such collapse can be expected under general circumstances for any operator which commutes with both the system Hamiltonian and the measured operator. Then, in section III we treat the particular case of a Heisenberg chain, finding that repeatedly measuring the z-component of the spin on the end of the chain causes it to oscillate with sharp Fourier peaks that correspond to energy gaps in the spectrum of the chain. We show that these oscillations arise as a result of the chain collapsing into a sector with a definite total z-component, making contact with our general proof. Finally, we comment on how these findings indicate that AO-QND measurement protocols may be of practical use in experiment and should not be overlooked.
@@ II Repeated Weak Measurement
Defining $\tau$ to be the time over which the pre-measurement is performed, the joint system-ancilla density matrix $\alpha$ becomes
@@
For completeness and since it will be used later, the expression for $\operatorname{Tr}_{\mu}(\operatorname{\mathbb{P}}_{\pm}\alpha_{k}(\tau))$ to second order is
$\operatorname{Tr}_{\mu}(\operatorname{\mathbb{P}}_{\pm}\alpha_{k}(\tau))$ $=\frac{\rho_{k}}{2}+\frac{i\tau}{2}[\rho_{k},H_{0}]+\frac{\tau^{2}}{2}(H_{0}\rho_{k}H_{0}-\frac{1}{2}\{H_{0}^{2},\rho_{k}\})\pm\frac{1}{2}\gamma\tau(\rho_{k}A+A\rho_{k})$ (7)
@@
II.1 Expectation Value of “Target” Observable
@@
Each $v_{nm}$ is a bounded super-martingale, i.e. each has the property that at all times $\mathbb{E}(v_{nm}(\rho_{k+1}))\leq v_{nm}(\rho_{k})$. Consequently, by the martingale convergence theorem it is guaranteed that each $v_{nm}$ converges to some stable value after a finite (but unknown) time [11]. We will now support the claim that the $v_{nm}$ are super-martingales, and give a condition for when they all converge to 00. It is evident that $v(\rho)$ is positive semi-definite (showing boundedness), and inspection reveals that $v(\rho)=0$ if and only if $\rho$ is collapsed into an eigensector of $B$. Thus, if we can show that $\mathbb{E}(v(\rho_{k+1})-v(\rho_{k}))\leq 0$ then $v(\rho)$ is a super-martingale and its convergence will tell us about collapse w.r.t. $B$.
@@
$\mathbb{E}(\Delta v_{mn})=-\frac{\gamma^{2}\tau^{2}}{2v_{mn}(\rho)^{3}}\left(\operatorname{Tr}(\operatorname{\mathbb{P}}_{n}\rho_{k})\operatorname{Tr}(\operatorname{\mathbb{P}}_{m}A\rho_{k})-\operatorname{Tr}(\operatorname{\mathbb{P}}_{m}\rho_{k})\operatorname{Tr}(\operatorname{\mathbb{P}}_{n}A\rho_{k})\right)^{2}.$ (11)
Since the term in parentheses is squared and we already know $v_{nm}\geq 0$, it is clear that $\mathbb{E}(\Delta v)\leq 0$ and so $v$ is a super-martingale which converges almost surely. There are then two remaining possibilities concerning the limit of $v$: either $v$ converges to 0 signifying that the state converges into a single eigensector of $B$, or $v$ converges to some other value. The latter case requires that an equilibrium be reached in which the term in parentheses is 00 at all times. Manipulating this term, we can rephrase this condition as being that the convergence stops at a step labeled $k_{0}$ when the condition
@@
is satisfied for all $n,m$ with $k\geq k_{0}$. Physically, this condition is that the expectation value of $A$ is at all times independent of any post-selection of an eigenvalue of $B$. To make contact with the QND case we again set $B=A$, giving the condition that $\langle A\rangle$ must not depend on any post-selection on $A$. This is clearly only met when the state occupies just one eigensector of $A$. So, we recover the QND result of guaranteed convergence to an eigenvalue. We therefore see that this result generalizes the second property of ordinary QND. This means that in our AO-QND protocol, measurement of the auxiliary observable $A$ really does cause the target observable $B$ to behave in a QND manner, thus effectively enabling a measurement-by-proxy of the target observable.
@@
Considering this condition more closely, we can see a direct link between information extraction and collapse. As seen in Eq. 5, the information extracted by the weak measurement is information about $\langle A\rangle$. Rephrasing the collapse-halting condition, we can say that the collapse stops when information about $\langle A\rangle$ ceases to be correlated with the value of $B$. So, we see that the weak measurement process indirectly extracts information about $B$, and that as long as it does so the system slowly collapses towards a $B$-eigensector. The collapse then stops either when an eigensector is reached and there is no new information about $B$ to be gained, or when eigensectors of $B$ can no longer be distinguished via $\langle A\rangle$ and so the weak measurement stops extracting new information about $B$. A similar result was obtained in [12], showing that an initially mixed state would gradually purify towards a subset of sectors as long as their signatures could be distinguished by the measurement. Here we have shown that this applies for arbitrary initial states, including pure states which are a superposition of contributions from multiple sectors.
@@
It is of further interest to examine what is required for a system to collapse only partially with respect to $B$. By partial collapse, we mean a situation in which the state does not become confined to a single $B$ eigensector, but rather reaches some superposition across multiple sectors which satisfies the collapse-halting condition of Eq. 12. The collapse-halting condition can be rephrased as requiring a state in which $\langle A\rangle$ is identical in two or more $B$-sectors for all $k>k_{0}$. That is, supposing we have some state with weight in two distinct $B$ eigensectors and $\lvert b_{n}\rangle,\lvert b_{m}\rangle$ are the projections of this state in those sectors, $\langle b_{n}\rvert A\lvert b_{n}\rangle_{k}=\langle b_{m}\rvert A\lvert b_{m}\rangle_{k}$. This necessarily requires $\langle A\rangle$ in each of the involved sectors to have the same time-dependence to maintain the balance. There are two sources of time-dependence acting on the system: the unitary evolution under which phases are introduced to energy eigenstates according to their eigenvalue, and the stochastic evolution introduced by the repeated measurements which adjust the weights of different $A$ eigenstates. Assuming that these two sources of time-dependence affect $\langle A\rangle$ independently over a single measurement cycle since one is deterministic and the other stochastic, both sources of time-dependence must therefore separately match across the involved sectors. To make this more concrete, consider the decomposition of $\lvert b_{n}\rangle$ into $A$ eigenstates: $\lvert b_{n}\rangle=\sum_{a}c_{n,a}\lvert n,a\rangle$, where $\lvert n,a\rangle$ represents a simultaneous eigenstate of $B$ and $A$ in the $n^{\textrm{th}}$ eigensector of $B$ with $A$ eigenvalue $a$. We can further rewrite this as $\lvert b_{n}\rangle=\sum_{a}\sum_{E}c_{n,a}c_{n,a,E}\lvert n,E\rangle$ by expanding the $\lvert n,a\rangle$ into a sum over simultaneous $B$ and energy eigenstates in the same sector. We can now argue that over some small timestep the Hamiltonian time-dependence acts only on $c_{n,a,E}$ by introducing a phase, while the time-dependence from weak measurement acts only on the $c_{n,a}$ by nudging these weights slightly in favor of some $a$ which is determined by the random measurement outcome. Thus, if we wish to find a non-collapsing state with weight in two sectors $n,m$, the decomposition in both sectors should be identical in the sense that $c_{n,a}=c_{m,a}$ and $c_{n,a,E}=c_{m,a,E}$. We have thus far not found any nontrivial examples of such a condition being met.
@@ III Application to Heisenberg Chain
Refer to caption(a)AmplitudeFrequency ($J/2\pi$)2$\leftrightarrow$1a1$\leftrightarrow$1b1$\leftrightarrow$1c2$\leftrightarrow$1b1$\leftrightarrow$1c2$\leftrightarrow$1c
@@
Figure 2: Behavior of the 4-site Heisenberg chain subjected to repeated weak measurement of $\sigma^{z}_{0}$. a) Fourier Transform of $\langle\sigma^{z}_{0}\rangle(t)$ showing sharp peaks corresponding to certain energy gaps in the spectrum of the chain. Labels on each peak show the contributing components, as identified in Table 1. b) Overlap of the chain’s state with sectors of total z-component shows a collapse into one sector following stochastic evolution. c) Distribution of expectation values of total z-component for a sample of 1,000 chains shown after certain numbers of measurements. It is clear that each individual run collapses into an eigensector of total z-component, while the ensemble average expectation value remains 0, demonstrating QND-like behavior.
@@
$H=-J\sum_{ij}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}$ (13)
We choose to label the spins along the chain from 00 to $N$ so that e.g. $\vec{\sigma}_{0}$ refers to the first spin on the chain, the ancilla. The operator we weakly measure is $\sigma^{z}_{0}$, the z-component of the first spin. To make the connection clear, this $\sigma^{z}_{0}$ operator corresponds to $A$ in the previous section and clearly does not commute with the Heisenberg Hamiltonian. The operator corresponding to $B$ will be the total z-component of the chain, which we refer to here as $m_{s}$. It is easy to verify that the required condition $[A,B]=[B,H]=0$ is met by this set of operators. The experimentally accessible information from this protocol will be the weak measurement record of $+1$ and $-1$ outcomes of the ancilla, but here we also explicitly track the system’s state over time in order to gain more insight into how it behaves.
@@
Of additional interest is the Fourier transform of $\langle\sigma^{z}_{0}\rangle$, which reveals oscillatory behavior in the z-component of the measured spin with several sharp peaks in frequency space. Examination of these peaks reveals that their spacing corresponds to certain spacings in the energy spectrum of the Heisenberg chain. More specifically, the oscillation frequencies correspond to energy spacings between eigenstates with the same total z-component (hence inhabiting the same $B$-sector) but with different energies and potentially different total spin, with allowed peaks following the selection rule $\Delta S=0,1$. In particular, the frequencies seen in Fig. 2a represent such spacings between eigenstates in the $m_{s}=1$ sector. That is, pairs of eigenstates with $m_{s}=1$ for both states and a difference of 0 or 1 in total spin contribute to the oscillation of the measured end spin. The fact that all possible peaks allowed under this rule are present indicates that following collapse into a total z-component sector, the effect of the measurement back-action causes the system state to explore all available states within the sector. It is yet to be determined whether this complete exploration of the sector is guaranteed for other systems, but there must be at least a partial exploration of the sector since by construction the measurement doesn’t commute with the Hamiltonian and so must perturb the system’s state to some extent. The finding of frequency peaks which obey this particular selection rule is specific to the Heisenberg chain but supports the finding of collapse into a single $m_{s}$ sector, as otherwise additional peaks would be present. These oscillations would also be seen in the experimentally accessible measurement record, though suppressed by noise. As such, our findings can be experimentally tested as long as a sufficient quantity of data is collected to overcome the statistical penalty of weak measurement. It is not clear at this time whether our protocol guarantees a similar experimental signature for arbitrary systems.
Total Spin Sector 2 1a 0a 1b 1c 0b
Total Spin 2 1 0 1 1 0
Energy -3 -1.83 -0.46 1 3.83 6.46
Table 1: Total spin sectors of the 4-site Heisenberg chain and their energies (J=1).